A function f(x) has the following form:

f(x)=kx 1<x<∞

and zero otherwise.

a)For what values of k is f(x) a pdf?

b)Find the CDF based on (a).

c)For what values of k does E(X) exist?


1
Expert's answer
2021-11-09T16:50:16-0500
"f(x) = \begin{cases}
 kx^{-(k+1)} &1<x<\infin \\
 0 & otherwise
\end{cases}"

a)

"\displaystyle\int_{-\infin}^{\infin}f(x)dx=1"


Then

"k\not=0"

"\displaystyle\int_{1}^{\infin}kx^{-(k+1)}dx=\lim\limits_{A\to\infin}\bigg[\dfrac{k}{-(k+1)+1}x^{-(k+1)+1}\bigg]\begin{matrix}
 A \\
 1
\end{matrix}"

"=-\lim\limits_{A\to\infin}A^{-k}+1=1"

"-\lim\limits_{A\to\infin}A^{-k}=0=>k>0"

Answer: "k>0"

b)


"F(x)=\displaystyle\int_{-\infin}^{x}f(y)dy"

"F(x)=0, x<1"

"F(x)=\displaystyle\int_{1}^{x} ky^{-(k+1)}dy"

"=\bigg[\dfrac{k}{-(k+1)+1}y^{-(k+1)+1}\bigg]\begin{matrix}
 x \\
 1
\end{matrix}=-x^{-k}+1"


Answer:

"F(x) = \begin{cases}
 0 &x<1 \\
 1-x^{-k} &1\leq x<\infin \\
\end{cases}, k>0"

c)


"E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{1}^{\infin}kx^{-(k+1)+1}dx"

"=\lim\limits_{A\to\infin}\bigg[\dfrac{k}{-(k+1)+2}x^{-(k+1)+2}\bigg]\begin{matrix}
 A \\
 1
\end{matrix}"

"=-\dfrac{k}{k-1}\lim\limits_{A\to\infin}A^{-k+1}+\dfrac{k}{k-1}, k>1"

If "k=1"


"E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{1}^{\infin}kx^{-1}dx"

"=\lim\limits_{A\to\infin}\big[k\ln(|x|)\big]\begin{matrix}
 A \\
 1
\end{matrix}=does\ not\ exist"


Answer:

 "E(X)" exists for "k>1."